Let x be a random variable that represents white blood cell count per cubic milliliter of whole blood. Assume that x has a distribution that is approximately normal, with mean μ = 6100 and estimated standard deviation σ = 2750. A test result of x < 3500 is an indication of leukopenia. This indicates bone marrow depression that may be the result of a viral infection.

(a) What is the probability that, on a single test, x is less than 3500? (Round your answer to four decimal places.)


(b) Suppose a doctor uses the average x for two tests taken about a week apart. What can we say about the probability distribution of x?

The probability distribution of x is approximately normal with μx = 6100 and σx = 1944.54.The probability distribution of x is approximately normal with μx = 6100 and σx = 2750.    The probability distribution of x is approximately normal with μx = 6100 and σx = 1375.00.The probability distribution of x is not normal.


What is the probability of x < 3500? (Round your answer to four decimal places.)


(c) Repeat part (b) for n = 3 tests taken a week apart. (Round your answer to four decimal places.)


(d) Compare your answers to parts (a), (b), and (c). How did the probabilities change as n increased?

The probabilities decreased as n increased.The probabilities stayed the same as n increased.    The probabilities increased as n increased.


If a person had x < 3500 based on three tests, what conclusion would you draw as a doctor or a nurse?

It would be an extremely rare event for a person to have two or three tests below 3,500 purely by chance. The person probably has leukopenia.It would be a common event for a person to have two or three tests below 3,500 purely by chance. The person probably does not have leukopenia.    It would be a common event for a person to have two or three tests below 3,500 purely by chance. The person probably has leukopenia.It would be an extremely rare event for a person to have two or three tests below 3,500 purely by chance. The person probably does not have leukopenia.
Answer #1

a)

P(Y < 3500) = P(Y - mean < 3500 - mean)
                  = P( (Y - mean)/SD < (3500 - mean)/SD
                  = P(Z < (3500 - mean)/SD)
                  = P(Z < (3500 - 6100)/2750)
                  = P(Z < -0.9455)
                  = 0.1722

b)

Mean = 6100, SD/root(N) = 1944.544

The probability distribution of x is approximately normal with μx = 6100 and σx = 1944.54.

P(Y < 3500) = P(Y - mean < 3500 - mean)
                  = P( (Y - mean)/(SD/root(N)) < (3500 - mean)/(SD/root(N))
                  = P(Z < (3500 - mean)/(SD/root(N)))
                  = P(Z < (3500 - 6100)/1944.544)
                  = P(Z < -1.3371)
                  = 0.0906

c)

Data given:
Mean = 6100, SD/root(N) = 1587.713, P(Y < 3500) = ?;

P(Y < 3500) = P(Y - mean < 3500 - mean)
                  = P( (Y - mean)/(SD/root(N)) < (3500 - mean)/(SD/root(N))
                  = P(Z < (3500 - mean)/(SD/root(N)))
                  = P(Z < (3500 - 6100)/1587.713)
                  = P(Z < -1.6376)
                  = 0.0507

d)

probabilities decreased as n increases.